OpenStudy (lena772):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x - 2y = -4 3x + y = 12 -2y - 4 2y - 4 2y + 4 -2y + 4

OpenStudy (ybarrap):

"isolate" means to put that variable by itself on one side of the equal sign. When you move things from one side of the equal sign to the other side, you must change its sign, positive numbers become negative and negative numbers become positive. So to "isolate" the x in the 1st equation, move it to the other side of the equal sign and change its sign. Next move the -4 to the other side of the equal sign and change its sign. This will be your equation.

OpenStudy (lena772):

-x - 2y = -4 -2y+4=x